OpenStudy (anonymous):

-65.376=7.2(9c+1) a. 1.12 c. –1.12 b. –0.9 d. –1.02

OpenStudy (jjcoolja12):

Let's solve your equation step-by-step. −65.376=7.2(9c+1) Step 1: Simplify both sides of the equation. −65.376=7.2(9c+1) Simplify: (Show steps) −65.376=64.8c+7.2 Step 2: Flip the equation. 64.8c+7.2=−65.376 Step 3: Subtract 7.2 from both sides. 64.8c+7.2−7.2=−65.376−7.2 64.8c=−72.576 Step 4: Divide both sides by 64.8. 64.8c 64.8 = −72.576 64.8 c=−1.12 Answer: c=−1.12

OpenStudy (anonymous):

thanks